Use technology to find the​ P-value for the hypothesis test described below.

The claim is that for 12 AM body​ temperatures, the mean is μ<98.6°F. The sample size is n=33 and the test statistic is t= - 2.515

P-value=??????
Answer #1

Solution:

Given that,

This is a left tailed test.

n = 33

df = n - 1 = 33 - 1 = 32

t = -2.515

Then,

p-value = 0.0086

The p-value is 0.0086
Answer #2

Using a t-distribution table or a calculator, the P-value for a two-tailed test with a test statistic of t = -2.515 and degrees of freedom (df) = 32 (n-1) is approximately 0.0179.


answered by: Hydra Master